OpenStudy (anonymous):

Express the ratio of the area of the larger circle to the area of the smaller circle in simplest radical form. Larger circle: radius of 2 + √3, smaller circle's radius: 2-√3.

OpenStudy (anonymous):



OpenStudy (anonymous):

The answer is supposed to be: 97 + 56√3

OpenStudy (dumbcow):

\[=\frac{\pi(2+\sqrt{3})^{2}}{\pi(2-\sqrt{3})^{2}} = \frac{7+4\sqrt{3}}{7-4\sqrt{3}}\]

OpenStudy (anonymous):

I approve

OpenStudy (dumbcow):

then you have to multiply by conjugate to get radical out of denominator

OpenStudy (anonymous):

Why is it 7 + 4√3?